Circular No. 5727

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

V346 NORMAE S. Molinari, R. Liseau and D. Lorenzetti, Istituto di Fisica dello Spazio

Interplanetario, Frascati; and J. Graham, Carnegie Institution, Washington, report: “We systematically observed this FU-Ori object (the exciting source for HH 57; cf. IAUC 3785, 3792, 3823) in the near-infrared during 1989-1992 with the European Southern Observatory’s 1-m telescope. From the literature, it appears that the object is steadily brightening, especially in the L and M bands, where it has undergone an increase of > 1.5 mag, along with a smaller fading in J. The last observation at ESO in Sept. 1992 gave J = 10.71 ± 0.03, H = 8.84 ± 0.01, K = 7.15 ± 0.01, L (3.8 μm)= 4.73 ± 0.02, M (4.6 μm)= 3.96 ± 0.05. A spectrum of the star taken with the Las Campanas 2.5-m telescope in June 1992 showed no significant changes from previous years, and we tentatively interpret the infrared brightening to evolution of the circumstellar dust. Further observations (especially spectroscopic) are strongly encouraged during the present season.”

COMET SHOEMAKER (1992y) Ephemeris continuation from the orbital elements on MPC 21533:

1993TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 3 2^{h}38.^{m}50 +60^{o}31.^{′}5 2.667 2.316 59.^{o}0 21.^{o}7 17.3
8 2 49.77 +61 52.7 2.691 2.319 57.9 21.4 17.3
13 3 02.42 +63 14.9 2.713 2.323 56.9 21.2 17.3
18 3 16.67 +64 37.4 2.733 2.329 56.2 21.0 17.4
23 3 32.75 +65 59.2 2.752 2.336 55.6 20.8 17.4
28 3 50.93 +67 19.0 2.769 2.344 55.2 20.6 17.4
May 3 4 11.51 +68 35.3 2.785 2.354 54.9 20.5 17.4
8 4 34.79 +69 46.1 2.801 2.365 54.7 20.4 17.5
13 5 01.0 +70 49.0 2.816 2.377 54.6 20.3 17.5
18 5 30.3 +71 41.4 2.830 2.391 54.6 20.2 17.5
23 6 02.5 +72 20.4 2.845 2.406 54.6 20.1 17.6
28 6 37.1 +72 43.2 2.861 2.421 54.7 20.0 17.6
June 2 7 13.1 +72 47.6 2.877 2.438 54.9 19.9 17.7
7 7 49.4 +72 32.4 2.894 2.457 55.0 19.8 17.7
12 8 24.6 +71 57.5 2.913 2.476 55.1 19.7 17.8
17 8 57.9 +71 04.4 2.933 2.496 55.2 19.5 17.8
22 9 28.48 +69 55.1 2.954 2.517 55.2 19.4 17.9

1993 March 27 Brian G. Marsden